OpenStudy (anonymous):

Find the sum of the following infinite geometric series if it exists. (2/5) + (12/25) + (72/125) +... A. 5/12 B. Does not exist C. 5/6 D. 6/5

OpenStudy (psymon):

\[\sum_{i=0}^{\infty} (\frac{ 2 }{ 5 })(\frac{ 6 }{ 5 })^{n}\] This is what your series appears to be. A geometric series in the form of ar^n diverges if r is greater than or equal to 1. Since r is 6/5, meaning greater than one, this series diverges and the sum would be infinite.

OpenStudy (anonymous):

So the answer is 6/5 since that is greater than one?

OpenStudy (anonymous):

@Psymon

OpenStudy (zzr0ck3r):

divide the 2nd term by the first term and you get r = (6/5) thus the sum does not exist at infinity because (6/5)>1

OpenStudy (zzr0ck3r):

\[\sum_{n=1}^{\infty}(\frac{6}{5})^n\]

OpenStudy (zzr0ck3r):

no need to start at 0

OpenStudy (psymon):

I've just been taught to start from 0 with geo series. I know it's not necessary, but its habit.

OpenStudy (anonymous):

Oh okay, I think I understand what you did there. If I had the following problem, how would I approach it? Find the sum of the following infinite geometric series if it exists. 1/2 + (-1/4) + 1/8 + (-1/16) +... It looks similar but I don't know.

OpenStudy (zzr0ck3r):

if you know its geometric, divide the second term by the first

OpenStudy (psymon):

Alright then. Dividing the 2nd term by the first.... \[\sum_{i=1}^{\infty}(\frac{ -1 }{ 2 })^{n}\] Since the absolute value of r is between 0 and 1, the sequence converges. If a geo series converges, it converges to the value of a/(1-r), meaning we would have: \[\frac{ 1 }{ 1-(\frac{ -1 }{ 2 })}\] = (2/3)

OpenStudy (zzr0ck3r):

\[\frac{\frac{1}{2}}{1-(\frac{-1}{2})}\]

OpenStudy (zzr0ck3r):

the first term is (1/2)

OpenStudy (psymon):

Yeah, dividing out the first term and ditching it just confused me, lol.

OpenStudy (zzr0ck3r):

\[\frac{first\space term}{1-ratio}\]

OpenStudy (psymon):

I'm just too used to keeping it there and starting from 0 x_x

OpenStudy (anonymous):

This is confusing lol. But the options I have are A. 1/3 B. 1/5 C. 1/12 D. 1/4

OpenStudy (anonymous):

Did we miss a step?

OpenStudy (psymon):

Nah, A is correct, I'm just trying to all of a sudden do something new and messing it up :P zz knows what he's doing.

OpenStudy (zzr0ck3r):

\[\frac{\frac{1}{2}}{1+\frac{1}{2}}=\frac{1/2}{\frac{3}{2}}=\frac{1}{2}*\frac{2}{3}=\frac{1}{3}\]

OpenStudy (zzr0ck3r):

just one little issue, but we all make that mistake...they normally start with 1

OpenStudy (zzr0ck3r):

give psy the medal he did most of it

OpenStudy (anonymous):

Oh okay, I forgot about bringing out 3/2 and making it 2/3. Thanks a lot for your help guys!

OpenStudy (psymon):

Yeah, when we learned it once upon a time ago, we always started from 0. Even my textbook starts it from 0.

OpenStudy (zzr0ck3r):

yeah

OpenStudy (psymon):

Oh well. Good to see something new :P